OpenStudy (anonymous):

Help please! -5(1-5x)+5(-8x-2)=-4x-8x solve for x

OpenStudy (anonymous):

-5+25x-40x-10=-12x -15x-15=-12x -3x=15 x=-5

OpenStudy (anonymous):

@ryan123345 @help123please.

OpenStudy (anonymous):

I just posted what you are supposed to do

OpenStudy (ryan123345):